The population standard deviation for the height of college basketball players is 3.1 inches. If we want to estimate 90% confidence interval for the population mean height of these players with a 1 margin of error, how many randomly selected players must be surveyed? (Round up your answer to nearest whole number)
Answer #1

Solution :

Given that,

standard deviation =s =  \sigma =3.1

Margin of error = E = 1

At 90% confidence level

\alpha = 1 - 90%  

\alpha = 1 - 0.90 =0.10

\alpha/2 = 0.05

Z\alpha/2 = Z0.05 = 1.645

sample size = n = [Z\alpha/2* \sigma / E] 2

n = ( 1.645* 3.1 /1 )2

n =26

Sample size = n =26